3x + 5y = 12 by the slope intercept method.
Answer by checkley71(8403)

3X+5Y=12 OR 5Y=-3X+12 OR Y=-3/5X+12/5 THUS THE SLOPE IS -3/5 & THE Y INTERCEPT IS 12/5
2 OF THE POINTS ARE (0,12/5) & (4,0)